OpenStudy (anonymous):

sin12x/sin7x find the limit

OpenStudy (anonymous):

with no doubt l'hopital not allowed

OpenStudy (kkutie7):

\[\frac{sin(12x)}{sin(7x)}\] here this is easier to look at

OpenStudy (anonymous):

\[\lim_{x\to 0}\frac{\sin(12x)}{\sin(7x)}\]

OpenStudy (anonymous):

Yes, but what is the correct answer? I keep getting 12/7. Not sure if this is right.

OpenStudy (anonymous):

what else could it be?

OpenStudy (anonymous):

you want the math teacher way?

OpenStudy (anonymous):

I want the helpful way..

OpenStudy (anonymous):

the snap way is l'hopital but i will guess that you did not get there yet the slow way is \[\lim_{x\to 0}\frac{\sin(ax)}{\sin(bx)}=\frac{\sin(ax)}{ax}\frac{bx}{\sin(bx)}\frac{ax}{bx}\]

OpenStudy (anonymous):

this requires knowing that \[\lim_{x\to 0}\frac{\sin(x)}{x}=1\] which you need

OpenStudy (anonymous):

Wow, never learned that, I know that, that is equal to one. thanks!

OpenStudy (anonymous):

yw

geerky42 (geerky42):

That... That is very good approach... wow I shed manly tear at the beauty of mathematics